OpenStudy (fanduekisses):

An object of mass m is at rest at the top of a smooth slope of height h and length L. The coefficient of kinetic friction between the object and the surface, mk, is small enough that the object will slide down the slope if given a very small push to get it started. Find an expression for the object’s speed at the bottom of the slope.

OpenStudy (fanduekisses):

wut?

OpenStudy (fanduekisses):

find acceleration and integrate it? idk how I'm supposed to solve this problem lol. This is a physics problem btw.

OpenStudy (fanduekisses):

@agent0smith

OpenStudy (agent0smith):

You don't need integration; acceleration will be constant. First thing, as always, is a free body diagram. Then work out the net force in the y and x directions.

OpenStudy (fanduekisses):

Like that?

OpenStudy (fanduekisses):

I asummed the slope to be on the x axis

OpenStudy (fanduekisses):

ohh ok

OpenStudy (fanduekisses):



OpenStudy (fanduekisses):

ok

OpenStudy (fanduekisses):

What am I solving for? acceleration?

OpenStudy (agent0smith):

Yes, using net force equations

OpenStudy (fanduekisses):

hehe \[F_{n}-mgcos \theta = m a \]

OpenStudy (fanduekisses):

y components

OpenStudy (agent0smith):

Yes, but a is zero for y

OpenStudy (fanduekisses):

\[F_{f}-mgsin \theta = ma\]

OpenStudy (fanduekisses):

look

OpenStudy (agent0smith):

Ff = mu*Fn

OpenStudy (fanduekisses):

\[a= \frac{ \mu_{k}F_{n}-mgsin \theta }{ m }\]

OpenStudy (agent0smith):

Use the y Fnet to solve for Fn.

OpenStudy (fanduekisses):

Why do I want to solve for Fn?

OpenStudy (agent0smith):

Because you have an Fn.

OpenStudy (fanduekisses):

\[F_{n}=\frac{ ma+mg \sin \theta }{ \mu_{k} }\]

OpenStudy (agent0smith):

"Yes, but a is zero for y"

OpenStudy (agent0smith):

Use the y-Fnet to solve for Fn, NOT the x-Fnet.

OpenStudy (fanduekisses):

ok \[F_n=mgcos \theta\]

OpenStudy (agent0smith):

Now you can plug it into your a equation and simplify.

OpenStudy (fanduekisses):

\[a= \frac{ \mu_{k}mgcos \theta - mgsin \theta }{ m }\]

OpenStudy (agent0smith):

Simplify. And keep in mind we don't know theta, but we have the height and length of the plane, so we can find cos and sine.

OpenStudy (fanduekisses):

\[a=g( \mu_k \cos \theta- \sin \theta) \]

OpenStudy (agent0smith):

Find cos and sine by drawing a triangle, and then once you have them, use \[\Large v_f^2 = v_i^2 + 2a x\]

OpenStudy (fanduekisses):

what is that?

OpenStudy (fanduekisses):

sin theta = h/x cos theta= L/ x

OpenStudy (agent0smith):



OpenStudy (fanduekisses):

ohh lol

OpenStudy (fanduekisses):

sin theta= h/ L cos theta = x/L x= L cos theta

OpenStudy (agent0smith):

Use pythag.

OpenStudy (fanduekisses):

\[\sqrt{L^2 - h^2}\]

OpenStudy (agent0smith):

Better.

OpenStudy (fanduekisses):

How do you go about representing speed?

OpenStudy (agent0smith):

The equation I gave you earlier, vf is final speed. We can assume initial is ~zero.

OpenStudy (agent0smith):

Plug in your big ugly expression for a. x would be the distance which is L

OpenStudy (fanduekisses):

\[v= \sqrt{2L[g(\mu_k \cos \theta-\sin \theta)}]\]

OpenStudy (fanduekisses):

:s

OpenStudy (agent0smith):

Good but you forgot sine and cosine

OpenStudy (fanduekisses):

lol so I just replace cos theta and sin theta with L/x and h/x respectively

OpenStudy (agent0smith):

There should be no x. That's why you used pythag...

OpenStudy (agent0smith):

And your sin theta is wrong. SOHCAHTOA.

OpenStudy (fanduekisses):

My bad I was looking at my first notes which were wrong yes so I meant sin theta= h/L and cos theta= sqrt(L^2-h^2)

OpenStudy (fanduekisses):

I'm so tired lol sorry

OpenStudy (agent0smith):

Now cos theta is wrong :P

OpenStudy (fanduekisses):

Ahhhhhhh :s

OpenStudy (fanduekisses):

cos theta= adj/hyp.... adj = sqrt(L^2-h^2)??

OpenStudy (fanduekisses):

ohhh typo lol I forgot to put it over L

OpenStudy (agent0smith):

I hate this question for not just saying " and the angle of the slope is theta". Makes the final expression so much worse looking.

OpenStudy (fanduekisses):

omg ikr

OpenStudy (agent0smith):

\[\large v= \sqrt{2Lg \left( \mu_k \frac{ \sqrt{L^2-h^2} }{ L }-\frac{ h }{ L }\right)} \]

OpenStudy (fanduekisses):

D:

OpenStudy (agent0smith):

Then cancel the common L \[\large v= \sqrt{2g \left( \mu_k \sqrt{L^2-h^2} - h \right)}\]who the F decided that this should be the expression instead of using theta.

OpenStudy (fanduekisses):

ooooohhh :}

OpenStudy (fanduekisses):

Ohhh well there is a part B to this and you're given actual values for height and length. look

OpenStudy (fanduekisses):

I can just plug in the values into the equation lol